# Trapping polar molecules by surface acoustic waves

### v1

@@ I Introduction
Polar molecules have many potential applications such as in quantum computation [1, 2, 3, 4, 5, 6] and simulations [7, 8, 9, 10]. The first step towards manipulating polar molecules is to decelerate and trap them for further operations [11]. It has been shown that polar molecules can be trapped by magnetic [12, 13, 14], optical [15, 16, 17, 18, 19] or electrical fields [20, 21, 22, 23, 24, 25, 19]. For examples, in the magneto electrostatic trap, polar molecules can be stably trapped by the magnetic quadrupole after deceleration [13]. By additionally applying both dc and ac electrical fields, the molecules can also be trapped via mechanical equilibrium [24, 26, 27]. Once successfully trapped, the internal rovibronic state, center of mass and interactions among different polar molecules can further be modulated via additional control fields [28, 29, 30], and variational efficient trapping approaches allow for more possible applications by controlling the trapped molecules [31, 32].
The interaction between electrical fields and polar molecules can be studied via the Stark effect that the electrical field can modulate the spectral properties of polar molecules [33, 34, 35, 36]. Based on this, polar molecules can be controlled or rotated by the electric field because of asymmetric structure and larger dipole moments, and this is different from the circumstance of neutral atoms [5]. Then the doublet splitting of polar molecules induced by the electric field can affect the Stark potential, which can further determine the mechanical potential, mechanical force, and mechanical motions of polar molecules [37, 38, 39, 40].
@@ II.1 Interaction between polar molecules and the electrical field by SAW
where $E_{\Lambda}=\bar{E}_{\Lambda}-kv$, $U_{j}=\pm\sqrt{\left(\frac{E_{\Lambda}}{2}\right)^{2}+\mu_{\rm eff}^{2}\left|\bar{E}_{z}^{(j)}\right|^{2}}$ representing the modulation of energy levels by the SAW provided electrical field, the effective dipole strength $\mu_{\rm eff}=\pm|\vec{\mu}|m\Omega/2J^{\prime}\left(J^{\prime}+1\right)$, which can be positive for the high-field seeking states or negative for the low-field seeking states [60, 37]. Obviously, $\mathbf{E}_{j}$ is independent of $\mathcal{E}_{x}$ in the RRF. In the following, we take $M=1$ for simplification.
@@
Figure 2: (a) The dependence of the energy levels of polar molecules on the electrical field with different quantum numbers, where the distance between the molecules and the lower SAW layer is $z=0.01$m, $E_{\Lambda}=0.4$GHz, $|\vec{\mu}|=0.167$Debye and the wave number $k=10$. (b) The energy level of polar molecules CO evaluated by $U_{j}\left(\bar{u}_{j},z\right)$ with $j=1,2$ representing the lower or upper SAW when $k=50$, $E_{\Lambda}=0.4$GHz, $|\vec{\mu}|=0.167$Debye, $J^{\prime}=m=\Omega=1$ [37].
@@ II.2 Trapping polar molecules into one-, two- and three-dimensional lattices
$dU_{j}/d\left|\bar{E}^{(j)}\right|=\mu_{\rm eff}\left|\bar{E}_{z}^{(j)}\right|/\sqrt{\left(E_{\Lambda}/2\right)^{2}+\left(\mu_{\rm eff}\left|\bar{E}_{z}^{(j)}\right|\right)^{2}}$, then
@@
$F_{z}^{(j)}(x,z,t)=\frac{\mu_{\rm eff}\bar{u}_{j}^{2}k^{3}e^{-2k\tilde{z}_{j}}}{\sqrt{\left(\frac{E_{\Lambda}}{2}\right)^{2}+\left(\mu_{\rm eff}\left|\bar{E}_{z}^{(j)}\right|\right)^{2}}}.$ (8b)
@@
$\vec{\nabla}\cdot\vec{F}^{(j)}(x,z,t)=-\frac{\mu_{\rm eff}\left[\left(\frac{\partial^{2}\phi}{\partial x^{2}}\right)^{2}+\left(\frac{\partial^{2}\phi}{\partial z^{2}}\right)^{2}+2\left(\frac{\partial^{2}\phi}{\partial z\partial x}\right)^{2}\right]}{\sqrt{\left(\frac{E_{\Lambda}}{2}\right)^{2}+\left(\mu_{\rm eff}\left|\vec{E}^{(j)}\right|\right)^{2}}}<0,$ (9)
where $\phi=\phi_{j}(x,z)$ is the electrical potential produced by trapping IDTs in the $j$th SAW layer, the inequality holds when $\mu_{\rm eff}>0$, and more details are given in Appendix C. Equation (9) indicates that the electrical field induced by SAW can trap the polar molecules if the condition of the equilibrium in Eq.(6) can be satisfied.
@@
When $\bar{u}_{1}$ and $\bar{u}_{2}$ are nonzero and there are no external forces, the polar molecules can be trapped by the electrical forces of IDTs between two piezoelectric materials, and the trapped location $z_{0}$ of molecules is determined by the equilibrium with $\bar{u}_{1}^{2}e^{-2kz_{0}}=\bar{u}_{2}^{2}e^{-2k(D-z_{0})}$, where $\bar{u}_{1}$ and $\bar{u}_{2}$ are proportional to the amplitudes of $u_{1}(t)$ and $u_{2}(t)$ in Fig. 1(a), respectively. Then we can obtain the trapping location $z_{0}=\ln\left(e^{2kD}\bar{u}_{1}^{2}/\bar{u}_{2}^{2}\right)/4k$, the relationship $e^{-kD}<\bar{u}_{1}/\bar{u}_{2}<e^{kD}$ always holds because $0<z_{0}<D$, and this agrees with the fact that larger $D$ induces more robust parameter settings of $\bar{u}_{1}$ and $\bar{u}_{2}$ for trapping the molecules. As shown in Fig. 3(a), the trapping locations are closer to the lower SAW layer when $\bar{u}_{1}>\bar{u}_{2}$ and are closer to the upper SAW layer when $\bar{u}_{2}>\bar{u}_{1}$. In this way, a single layer of polar molecules can be trapped at arbitrary locations of the height $z_{0}$. Besides, here the potential applied on the polar molecules is the joint potential generated by the trapping IDTs of the upper and lower SAW layers. Then Eq. (92) in Appendix C reveals that the molecules can be stably trapped at the equilibrium positions.
@@
$=\frac{\mu_{\rm eff}M^{2}\bar{u}_{1}^{2}k^{3}e^{-2kz_{0}}}{\sqrt{\left(E_{\Lambda}/2\right)^{2}+\left(\mu_{eff}|\bar{E}^{(1)}|\right)^{2}}}$
$\approx M\bar{u}_{1}k^{2}e^{-kz_{0}}$
@@
with $f_{1}=M\bar{u}_{1}$. Then $\tilde{F}_{z}(x_{0},z_{0},t)$ can be used to cancel $F_{z}^{(1)}(x_{0},z_{0},t)$ based on the approximation $E_{\Lambda}\ll\left|\bar{E}^{(1)}(x,z,t)\right|$ [47].
@@
Further for the simulations in Fig. 3(b), the red curve denoting external electric force has two intersection points $\mathrm{P}$ and $\mathrm{Q}$ with the SAW induced electrical force. For the lower intersection point $\mathrm{P}$ at the height $z_{p}$, the combined force is downward when $z<z_{p}$ and upward when $z>z_{p}$, thus the trapped polar molecules can easily escape from the trapped site at $z_{p}$. However, for the upper intersection point $\mathrm{Q}$ at the height $z_{q}$, the combined force is upward when $z<z_{q}$ and downward when $z>z_{q}$, and this provides a resilience along $z$ direction to make sure that the polar molecules are stably trapped at $z_{q}$. This is why in Fig.3(c), the black-dot curve can trap the polar molecules at the desired layers more stably then the green-dash curve. Although the black-dot curve can simultaneously introduce unstable trapped intersections, while the number of trapped polar molecules is much less than that by the stable intersections.
Once the polar molecules are stably trapped by the electrical force, the interactions among molecules are affected by their locations and external electrical potentials. In the following, we discuss three different cases: the first is that the polar molecules are trapped as a three-dimensional array, the lattices sites in each layer are remotely separated in the horizontal direction and the horizontal hoppings are not considered, the second is that a single layer of polar molecules are trapped and the horizontal hoppings are considered, and the third is that the polar molecules are trapped by one SAW layer (i.e., the upper layer) and the external force in lattices, but their quantum dynamics are also affected by the electrical potential induced by the other SAW layer, then the lattice dynamics can be influenced by the lower SAW.
@@ III Three-dimensional trapped polar molecule array with multiple layers
$H_{L}=\sum_{l=1}^{L}\left[\frac{P_{l}^{2}}{2m_{0}}+\frac{m_{0}\omega^{2}r_{l}^{2}}{2}\right]+\frac{1}{2}\sum_{q\neq l}^{L}V_{|q-l|}\left(|z_{i}-z_{l}|\right),$ (12)
@@
as a generalization of the formula in Ref. [63], the parameter $R_{l}$ is determined by the trapped locations of different polar molecule layers. For the multi-layer trapped molecules, the binding energy represents the required energy to separate the layers as $E_{b}=\sum_{q\neq l}E_{b}^{ql}\approx-\sum_{q\neq l}\exp(-\alpha(z_{q}-z_{l})^{2})$ [64].
@@ IV Spatial distribution of trapped molecules
Applying the Laplace transform to Eq. (15), then we have
@@
Let us assume initially $p_{1}(0)=1$ and $p_{n}(0)=0$ for $n=2,3,\cdots,N$, that is, only molecules at the first lattice are initially trapped. Then Eq. (16) with $n=1$ reads [65]
@@
where $P_{j^{\prime}s}\approx\frac{1}{is-U_{0}}v_{j^{\prime}1}P_{1}(s)$ has been used. Then
@@
$+\frac{1}{(is-U_{0})^{3}}\sum_{j=2}^{N}v_{1j}\left[\sum_{j^{\prime}=2}^{j-1}v_{jj^{\prime}}v_{j^{\prime}1}+\sum_{j+1}^{N}v_{jj^{\prime}}v_{j^{\prime}1}\right]P_{1}(s).$
Denoting $\alpha=\sum_{j=2}^{N}v_{1j}^{2}$ and $\beta=\sum_{j=2}^{N}v_{1j}\left[\sum_{j^{\prime}=2}^{j-1}v_{jj^{\prime}}v_{j^{\prime}1}+\sum_{j+1}^{N}v_{jj^{\prime}}v_{j^{\prime}1}\right]$, then we have
@@
Then the trapped molecule array in Fig. 3(a) can be regarded as a lattice network with infinite number of sites and the spatial distribution of trapped localizations converges with the increase of $N$, as compared in Figs. 4(a) and (b).
Figure 4: (a) Comparison of the probability of trapping at different localized sites with two different single-layer trapping methods. $U_{0}=2$ in both of simulations, $v_{nj}=0.4\left|x_{n}-x_{j}\right|^{-3}$, $x_{n}=0.1n$, $N=5$ for discrete trapping, and the discrete trapping converges to the continuous case when $N=500$. (b) The blue-dashed-circle evaluated by left-y axis represents the steady value of $P_{1}$ between $\max_{n}\left(P_{n}\right)$ and $\min_{n}\left(P_{n}\right)$ with the same parameters as (a), and decreases with the increase of lattice sites. The diamond is for the cooperative shielding that $|\psi(0)\rangle=1/\sqrt{N}\sum_{n=1}^{N}|n\rangle$, the asterisk is for that $|\psi(0)\rangle=1/\sqrt{5}\sum_{n=1}^{5}|n\rangle$, and both of the two simulations are evaluated by right-y axis. (c) Comparison of cooperative shielding with $|c_{n}(T)|$ when $|\psi(0)\rangle=1/\sqrt{N}\sum_{n=1}^{N}|n\rangle$ between $N=100$ (almost continuous trapping) and $N=10$ (inset, discrete trapping). (d) Comparison when $|\psi(0)\rangle=1/\sqrt{5}\sum_{n=1}^{5}|n\rangle$. In (b-e), $V=1$ and $U_{0}=0.1$.
@@ IV.1 Comparison of diffusion based on the single-layer trapped polar molecules
Generalized from the case in Eq. (15), when the polar molecules are trapped by two SAW layers as shown in Fig. 1(a) without the external electric force, the molecules can be trapped at arbitrary locations along the $x$ direction, then the spatial distribution of the molecules is governed by

### v4

@@ I Introduction
Polar molecules have many potential applications such as in quantum computation [1, 2, 3, 4, 5, 6] and simulations [7, 8, 9, 10]. The first step towards manipulating polar molecules is to decelerate and trap them for further operations [11]. It has been shown that polar molecules can be trapped by magnetic [12, 13, 14], optical [15, 16, 17, 18, 19] or electrical fields [20, 21, 22, 23, 24, 25, 19]. For examples, in the magneto electrostatic trap, polar molecules can be stably trapped by the magnetic quadrupole after deceleration [13]. By additionally applying both dc and ac electrical fields, the molecules can also be trapped via mechanical equilibrium [24, 26, 27]. Once successfully trapped, the internal rovibronic state, center of mass and interactions among different polar molecules can further be modulated via additional control fields [28, 29, 30]. Variational efficient trapping approaches allow for more possible applications by controlling the trapped molecules [7, 31].
The interaction between electrical fields and polar molecules can be studied via the Stark effect that the electrical field can modulate the spectral properties of polar molecules [32, 33, 34, 35]. Based on this, polar molecules can be controlled or rotated by the electric field because of asymmetric structure and larger dipole moments, and this is different from the circumstance of neutral atoms [5]. The doublet splitting of polar molecules induced by the electric field can affect the Stark potential, which can further determine the mechanical potential, mechanical force, and mechanical motions of polar molecules [36, 37, 38, 39].
@@ II.1 Interaction between polar molecules and the electrical field by SAW
where $E_{\Lambda}=\bar{E}_{\Lambda}-kv$, $U_{j}=\pm\sqrt{\left(E_{\Lambda}/2\right)^{2}+\mu_{\rm eff}^{2}\left|\bar{E}_{z}^{(j)}\right|^{2}}$ representing the modulation of energy levels by the SAW provided electrical field, the effective dipole strength $\mu_{\rm eff}=\pm|\vec{\mu}|m\Omega/2J^{\prime}\left(J^{\prime}+1\right)$, which can be positive for the high-field seeking states or negative for the low-field seeking states [59, 36]. Obviously, $\mathbf{E}_{j}$ is independent of $\mathcal{E}_{x}$ in the RRF. In the following, we take $M=1$ for simplification.
@@
Figure 2: (a) The dependence of the energy levels of polar molecules on the electrical field with different quantum numbers, where the distance between the molecules and the lower SAW layer is $z=0.01$m, $E_{\Lambda}=0.4$GHz, $|\vec{\mu}|=0.167$Debye and the wave number $k=10$. (b) The energy level of polar molecules CO evaluated by $\max\left(U_{j}\left(\bar{u}_{j},z\right)\right)$ with $j=1,2$ representing the lower or upper SAW when $k=50$, $E_{\Lambda}=0.4$GHz, $|\vec{\mu}|=0.167$Debye, $J^{\prime}=m=\Omega=1$ [36].
@@ II.2 Trapping polar molecules into one-, two- and three-dimensional lattices
$dU_{j}/d\left|\bar{E}^{(j)}_{z}\right|=\mu_{\rm eff}^{2}\left|\bar{E}_{z}^{(j)}\right|/\sqrt{\left(E_{\Lambda}/2\right)^{2}+\left(\mu_{\rm eff}\left|\bar{E}_{z}^{(j)}\right|\right)^{2}}$, then
@@
$F_{z}^{(j)}(x,z,t)=\frac{\mu_{\rm eff}^{2}\bar{u}_{j}^{2}k^{3}e^{-2k\tilde{z}_{j}}}{\sqrt{\left(\frac{E_{\Lambda}}{2}\right)^{2}+\left(\mu_{\rm eff}\left|\bar{E}_{z}^{(j)}\right|\right)^{2}}}.$ (8b)
@@
$\vec{\nabla}\cdot\vec{F}^{(j)}(x,z,t)=-\frac{\mu_{\rm eff}^{2}\left[\left(\frac{\partial^{2}\phi}{\partial x^{2}}\right)^{2}+\left(\frac{\partial^{2}\phi}{\partial z^{2}}\right)^{2}+2\left(\frac{\partial^{2}\phi}{\partial z\partial x}\right)^{2}\right]}{\sqrt{\left(\frac{E_{\Lambda}}{2}\right)^{2}+\left(\mu_{\rm eff}\left|\vec{E}^{(j)}\right|\right)^{2}}}<0,$ (9)
where $\phi=\phi_{j}(x,z)$ is the electrical potential produced by trapping IDTs in the $j$th SAW layer and more details are given in Appendix C. Equation (9) indicates that the electrical field induced by SAW can trap the polar molecules if the condition of the equilibrium in Eq.(6) can be satisfied.
@@
When $\bar{u}_{1}$ and $\bar{u}_{2}$ are nonzero and there are no external forces, the polar molecules can be trapped by the electrical forces of IDTs between two piezoelectric materials, and the trapped location $z_{0}$ of molecules is determined by the equilibrium with $\bar{u}_{1}^{2}e^{-2kz_{0}}=\bar{u}_{2}^{2}e^{-2k(D-z_{0})}$, where $\bar{u}_{1}$ and $\bar{u}_{2}$ are proportional to the amplitudes of $u_{1}(t)$ and $u_{2}(t)$ in Fig. 1(a), respectively. Then we can obtain the trapping location $z_{0}=\ln\left(e^{kD}\bar{u}_{1}/\bar{u}_{2}\right)/2k$, the relationship $e^{-kD}<\bar{u}_{1}/\bar{u}_{2}<e^{kD}$ always holds because $0<z_{0}<D$, and this agrees with the fact that larger $D$ induces more robust parameter settings of $\bar{u}_{1}$ and $\bar{u}_{2}$ for trapping the molecules. As shown in Fig. 3(a), the trapping locations are closer to the lower SAW layer when $\bar{u}_{1}>\bar{u}_{2}$ and are closer to the upper SAW layer when $\bar{u}_{2}>\bar{u}_{1}$. In this way, a single layer of polar molecules can be trapped at arbitrary locations of the height $z_{0}$. Besides, here the potential applied on the polar molecules is the joint potential generated by the trapping IDTs of the upper and lower SAW layers. Then Eq. (92) in Appendix C reveals that the molecules can be stably trapped at the equilibrium positions.
@@
$=\frac{\mu_{\rm eff}^{2}M^{2}\bar{u}_{1}^{2}k^{3}e^{-2kz_{0}}}{\sqrt{\left(E_{\Lambda}/2\right)^{2}+\left(\mu_{\rm eff}\left|\bar{E}_{z}^{(1)}\right|\right)^{2}}}$
@@
with $f_{1}\approx\mu_{\rm eff}M\bar{u}_{1}$. $\tilde{F}_{z}(x_{0},z_{0},t)$ can be used to cancel $F_{z}^{(1)}(x_{0},z_{0},t)$ based on the approximation $E_{\Lambda}\ll\left|\bar{E}^{(1)}(x,z,t)\right|$ [46].
@@
Further for the simulations in Fig. 3(b), the red curve denoting external electric force has two intersection points $\mathrm{P}$ and $\mathrm{Q}$ with the SAW induced electrical force. For the lower intersection point $\mathrm{P}$ at the height $z_{p}$, the combined force is downward when $z<z_{p}$ and upward when $z>z_{p}$, thus the trapped polar molecules can easily escape from the trapped site at $z_{p}$. However, for the upper intersection point $\mathrm{Q}$ at the height $z_{q}$, the combined force is upward when $z<z_{q}$ and downward when $z>z_{q}$, and this provides a resilience along $z$ direction to make sure that the polar molecules are stably trapped at $z_{q}$. This is why in Fig.3(c), the black-dot curve can trap the polar molecules at the desired layers more stably than the green-dash curve. Although the black-dot curve can simultaneously introduce unstable trapped intersections, while the number of trapped polar molecules is much less than that by the stable intersections.
Once the polar molecules are stably trapped by the electrical force, the interactions among molecules are affected by their locations and external electrical potentials. In the following, we discuss three different cases: the first is that the polar molecules are trapped as a three-dimensional array, the lattice sites in each layer are remotely separated in the horizontal direction and the horizontal hoppings are not considered, the second is that a single layer of polar molecules are trapped and the horizontal hoppings are considered, and the third is that the polar molecules are trapped by one SAW layer (i.e., the upper layer) and the external force in lattices, but their quantum dynamics are also affected by the electrical potential induced by the other SAW layer, so that the lattice dynamics can be influenced by the lower SAW.
@@ III Three-dimensional trapped polar molecule array with multiple layers
$H_{L}=\sum_{l=1}^{L}\left(\frac{P_{l}^{2}}{2m_{0}}+\frac{m_{0}\omega^{2}r_{l}^{2}}{2}\right)+\frac{1}{2}\sum_{q\neq l}^{L}V_{|q-l|}\left(|z_{i}-z_{l}|\right),$ (12)
@@
as a generalization of the formula in Ref. [62], and the parameter $R_{l}$ is determined by the trapped locations of different polar molecule layers. For the multi-layer trapped molecules, the binding energy representing the required energy to separate the layers reads $E_{b}=\sum_{q\neq l}E_{b}^{ql}\approx-\sum_{q\neq l}\exp(-\alpha(z_{q}-z_{l})^{2})$ [63].
@@ IV Spatial distribution of trapped molecules
Applying the Laplace transform to Eq. (15), we have
@@
Let us assume initially $p_{1}(0)=1$ and $p_{n}(0)=0$ for $n=2,3,\cdots,N$, that is, only molecules at the first lattice are initially trapped. Equation (16) with $n=1$ reads [64]
@@
where $P_{j^{\prime}}(s)\approx\frac{1}{is-U_{0}}v_{j^{\prime}1}P_{1}(s)$ has been used. Then
@@
$+\frac{1}{(is-U_{0})^{3}}\sum_{j=2}^{N}v_{1j}\left(\sum_{j^{\prime}=2}^{j-1}v_{jj^{\prime}}v_{j^{\prime}1}+\sum_{j+1}^{N}v_{jj^{\prime}}v_{j^{\prime}1}\right)P_{1}(s).$
Denoting $\alpha=\sum_{j=2}^{N}v_{1j}^{2}$ and $\beta=\sum_{j=2}^{N}v_{1j}\left(\sum_{j^{\prime}=2}^{j-1}v_{jj^{\prime}}v_{j^{\prime}1}+\sum_{j+1}^{N}v_{jj^{\prime}}v_{j^{\prime}1}\right)$, then we have
@@
The trapped molecule array in Fig. 3(a) can be regarded as a lattice network with infinite number of sites and the spatial distribution of trapped localizations converges with the increase of $N$, as compared in Figs. 4(a) and (b).
Figure 4: (a) Comparison of the probability of trapping at different localized sites with two different single-layer trapping methods. $U_{0}=2$ in both of simulations, $v_{nj}=0.4\left|x_{n}-x_{j}\right|^{-3}$, $x_{n}=0.1n$, $N=5$ for discrete trapping, and the discrete trapping converges to the continuous case when $N=500$. (b) The blue-dashed-circle evaluated by left-y axis represents the steady value of $P_{1}$ between $\max_{n}\left(P_{n}\right)$ and $\min_{n}\left(P_{n}\right)$ with the same parameters as (a), and decreases with the increase of lattice sites. The diamond is for the cooperative shielding that $|\psi(0)\rangle=1/\sqrt{N}\sum_{n=1}^{N}|n\rangle$, the asterisk is for that $|\psi(0)\rangle=1/\sqrt{5}\sum_{n=1}^{5}|n\rangle$, and both of the two simulations are evaluated by right-y axis. (c) Comparison of cooperative shielding with $|c_{n}(T)|$ when $|\psi(0)\rangle=1/\sqrt{N}\sum_{n=1}^{N}|n\rangle$ between $N=100$ (almost continuous trapping) and $N=10$ (inset, discrete trapping). (d) Comparison when $|\psi(0)\rangle=1/\sqrt{5}\sum_{n=1}^{5}|n\rangle$. In (b-d), $V=1$ and $U_{0}=0.1$.
@@ IV.1 Comparison of diffusion based on the single-layer trapped polar molecules
Generalized from the case in Eq. (15), when the polar molecules are trapped by two SAW layers as shown in Fig. 1(a) without the external electric force, the molecules can be trapped at arbitrary locations along the $x$ direction, the spatial distribution of the molecules is governed by